OpenStudy (anonymous):

Find the x and y coordinates. y=x^3 - 4x y= -(x + 2) Please don't tell me the answer. i'd like to know how to work the problem. thanks

OpenStudy (valpey):

These are two different curves or lines in the xy plane. They intersect at specific point(s) which will be places where the values for x and y solve both equations. You can start by setting the two equations equal to each other to find x values where this could be true.

OpenStudy (anonymous):

i can't. i have to use substitution or elimination or some other method and i've tried...i'm stuck..

OpenStudy (anonymous):

they are two different lines and they are supposed to intersect at what looks like two points maybe..but idk where to go from here

OpenStudy (valpey):

\[y=x^3-4x\]\[y=-(x+2)\]\[x^3-4x=-(x+2)\]\[x^3-4x+x+2=0\]Can you factor this?

OpenStudy (anonymous):

yeah thats basically what i did...its the 4x + x thats throwing me off because that won't factor

OpenStudy (valpey):

-4x+x = -3x

OpenStudy (valpey):

You are missing an x^2 term. But this does factor quite neatly.

OpenStudy (anonymous):

okay....i already know all that..lol...sorry my brain is fried

OpenStudy (anonymous):

okay i found it..careless mistake..i was getting something way different

OpenStudy (valpey):

What I like to do is first think, are there any obvious positive solutions? Are there any obvious negative solutions? Start with the most simple ones (things that divide 2, for example)

OpenStudy (anonymous):

i can't think ahead..i just work the problems..i know how to fix a problem not do it in my head..but i'm still getting -x^3 + 4x -2

OpenStudy (anonymous):

=x

OpenStudy (anonymous):

how does the x go over like that?i'm so confused

OpenStudy (valpey):

Oh, you are not following: If y = Equation 1 and y = Equation 2 then Equation 1 = Equation 2

OpenStudy (anonymous):

thats how i set up...then i moved the 2 over then the negative..

OpenStudy (valpey):

You should have \[0 = -x^3 + 3x-2\]or\[x^3-3x+2=0\] Did you do this?: \[y=x^3−4x\]\[y=−(x+2)\]\[x^3−4x=−(x+2)\]\[x^3−4x+x+2=0\]\[x^3−3x+2=0\]

OpenStudy (anonymous):

yeah...i had plus 4x instead of -4x..woops

OpenStudy (anonymous):

do i use synthetic division now?

OpenStudy (valpey):

Okay, so can you factor using the obvious solutions?

OpenStudy (valpey):

(the obvious solutions to try are going to be the plus or minus divisors of 2)

OpenStudy (anonymous):

i can't do that with a cubed x..i thought that only worked with squares

OpenStudy (valpey):

Just plug in 1, -1, 2,and -2 into the equation and see if any of them work.

OpenStudy (anonymous):

ummm...i think i can handle it from here..thank you...